# Time Series Anomaly Detection in the Frequency Domain with Statistical Reliability

### v1

@@ Abstract
Effective anomaly detection in complex systems requires identifying change points (CPs) in the frequency domain, as abnormalities often arise across multiple frequencies. This paper extends recent advancements in statistically significant CP detection, based on Selective Inference (SI), to the frequency domain. The proposed SI method quantifies the statistical significance of detected CPs in the frequency domain using $p$-values, ensuring that the detected changes reflect genuine structural shifts in the target system. We address two major technical challenges to achieve this. First, we extend the existing SI framework to the frequency domain by appropriately utilizing the properties of discrete Fourier transform (DFT). Second, we develop an SI method that provides valid $p$-values for CPs where changes occur across multiple frequencies. Experimental results demonstrate that the proposed method reliably identifies genuine CPs with strong statistical guarantees, enabling more accurate root-cause analysis in the frequency domain of complex systems.
@@ 1 Introduction
To detect failures in complex systems, it is crucial to accurately identify the frequencies where the anomalies occur. By pinpointing these specific frequencies, the root causes of faults in the systems can be identified. In this paper, we address the change point (CP) detection problem in the frequency domain and propose a method that can identify statistically significant changes in specific frequencies. Quantifying statistical significance in CP detection ensures that the detected changes in the frequency domain reflect genuine structural alterations rather than random noise. Measures such as $p$-values provide a quantitative framework to differentiate real changes from spurious ones, effectively mitigating the risk of incorrect decisions in data-driven systems.
@@
One of the difficulties in frequency-domain CP detection is that changes often appear across multiple frequencies due to shared underlying phenomena influencing broad signal characteristics. Figure 1 shows an example of frequency-domain CP detection problems along with the statistically significant CPs identified using our proposed SI method. Panel (a) shows a time series transformed into signals across five frequencies. Panel (b) presents the CP detection results for each individual frequency $d\in\{d_{0},d_{1},d_{2},d_{3},d_{4}\}$ 111 Note that while the “amplitude” spectra are presented in the figures of this paper to visualize the temporal variations of spectral sequences and the means for each segment, the “complex” spectra are utilized in the actual CP detection and the hypothesis testing (see Section 2 and the subsequent sections for details). . Panel (c) integrates CPs across multiple frequencies, detecting three changes A, B, and C (at CP A, changes occur in frequencies $d_{1}$ and $d_{2}$; at CP B, in $d_{2}$ and $d_{3}$; and at CP C, only in $d_{0}$). In table (d), the obtained $p$-values for changes A, B, and C using the proposed method are 0.008, 0.006, and 0.752 (denoted as selective $p$-values), respectively, indicating that at a significance level of $\frac{0.05}{3}\approx 0.0167$ decided by Bonferroni correction, changes A and B are statistically significant CPs222 In Figure 1, the values labeled as naive $p$-values are inappropriate because they fail to account for double-dipping, leading to excessively small values and an inflated rate of false positive findings (see Section 4 for details). .
@@
To perform statistically significant CP detection, as illustrated in Figure 1, two key technical challenges must be addressed. The first challenge is to extend the SI framework to the frequency domain. We tackle this by formulating the test statistic and the conditioning on the hypothesis selection that accurately account for the properties of discrete Fourier transform (DFT). The second challenge involves quantifying the statistical significance of CPs shared across multiple frequencies. Addressing this requires solving a combinatorial optimization problem, which is computationally infeasible to solve globally optimally, necessitating a heuristic approach to obtain the approximate solution. In this study, we employ a method based on simulated annealing (Kirkpatrick et al., 1983; Černỳ, 1985) and implement the SI framework to appropriately quantify the statistical significance of the approximate solution derived from the heuristic algorithm.
The rest of the paper is organized as follows. Section 2 formulates the problem. Section 3 explains the heuristic algorithm used to select CP candidates in the first stage of the proposed method. Section 4 describes the method for quantifying the statistical significance of CP candidates using the SI framework in the second stage. In Section 5, we demonstrate the effectiveness of our proposed method through comprehensive numerical experiments using both synthetic and real-world data. For reproducibility, our implementation is available at https://github.com/Takeuchi-Lab-SI-Group/si_for_frequency-domain_anomaly_detection. Finally, Section 6 concludes the paper. We note that the selection of CP candidates in stage 1 (Section 3) does not involve any particularly technical contributions. Our primary contribution lies in stage 2 (Section 4), where the statistical significance of the approximate solution is evaluated using the SI framework.
Related Work. The CP detection problem has long been studied with various applications in a variety of fields, such as finance (Fryzlewicz & Subba Rao, 2014; Pepelyshev & Polunchenko, 2017), bioinformatics (Chen & Wang, 2008; Muggeo & Adelfio, 2011; Pierre-Jean et al., 2015), climatology (Reeves et al., 2007; Beaulieu et al., 2012), and machine monitoring (Lu et al., 2017; 2018). In the statistics and machine learning communities, various methods have been proposed for identifying multiple CPs from univariate sequences. The most straightforward approach is repeatedly applying single CP detection algorithms, such as binary segmentation (Scott & Knott, 1974). Examples of such approaches include circular binary segmentation (Olshen et al., 2004) and wild binary segmentation (Fryzlewicz, 2014). Another line of research has proposed numerous approaches based on penalized likelihood, such as Segment Neighbourhood (Auger & Lawrence, 1989), Optimal Partitioning (Jackson et al., 2005), PELT (Killick et al., 2012), and FPOP (Maidstone et al., 2017). In these studies, dynamic programming is employed to solve the penalized likelihood minimization problem. While most CP detection studies focus on the time domain, a few studies have targeted the frequency domain, such as Adak (1998), Last & Shumway (2008), and Preuss et al. (2015).
@@
SI was initially introduced as a method of statistical inference for feature selection in linear models (Taylor & Tibshirani, 2015; Fithian et al., 2015) and later extended to various feature selection algorithms, including marginal screening (Lee & Taylor, 2014), stepwise feature selection (Tibshirani et al., 2016), and Lasso (Lee et al., 2016). The core concept of SI is to derive the exact null distribution of the test statistic conditional on the hypothesis selection event, thereby enabling valid statistical inference with controlled type I error rate. Recently, significant attention has been given to applying SI to more complex supervised learning algorithms, such as kernel models (Yamada et al., 2018), boosting (Rügamer & Greven, 2020), tree-structured models (Neufeld et al., 2022), and neural networks (Duy et al., 2022; Miwa et al., 2023; Shiraishi et al., 2024b). Furthermore, SI has proven valuable for unsupervised learning tasks, including clustering (Lee et al., 2015; Chen & Witten, 2023; Gao et al., 2024), outlier detection (Chen & Bien, 2020; Tsukurimichi et al., 2022), domain adaptation (Duy et al., 2024), and segmentation (Tanizaki et al., 2020; Duy et al., 2022). SI was first used for CP detection problem in (Hyun et al., 2018), which focused on Fused Lasso algorithm. Since then, the framework has been explored in various CP detection algorithms, including the CUSUM-based method (Umezu & Takeuchi, 2017), binary segmentation and its variants (Hyun et al., 2021), dynamic programming (Duy et al., 2020), and other related problems (Sugiyama et al., 2021; Jewell et al., 2022; Carrington & Fearnhead, 2024; Shiraishi et al., 2024a). Existing studies on CP detection have focused on the time domain, making this the first to offer valid statistical inferences for frequency-domain anomalies using the SI framework.
@@ 2.2 Statistically Significant CP Detection in the Frequency Domain
As discussed in Section 1, we have prior knowledge that simultaneous changes occur across multiple distinct frequencies. In Section 3, we introduce a heuristic algorithm that generates CP candidates by incorporating this prior knowledge. Subsequently, in Section 4, we present an SI framework to quantify the statistical significance of each CP candidate in the form of $p$-values. Finally, we select the candidates with $p$-values smaller than the user-specified significance level (e.g., 0.05 or 0.01) as our final CPs. The flow of CP detection in the frequency domain is illustrated in Figure 2.
Figure 2: Schematic illustration of CP detection process in the frequency domain. In this example, at CP 28, changes occur in frequencies $d_{1}$ and $d_{2}$; at CP 70, in $d_{2}$ and $d_{3}$; and at CP 97, only in $d_{0}$. Our proposed SI method in Section 4 can provide valid $p$-values for each CP.
@@ 3 Heuristic Algorithm for CP Candidate Selection in Multiple Frequencies
As discussed in Section 1, the types of anomalies occurring in sensor signals can be systematically determined by simultaneously detecting multiple frequency anomalies, such as the harmonics and sidebands of the characteristic frequencies. Thus, we formulate CP candidate selection as an optimization problem that aims to not only estimate the optimal number and location of CPs for each frequency, but also reduce the total number of change locations across all frequencies by aligning the positions.
@@ 3.1 Objective Function for CP Candidate Selection
Let $K^{(d)}$ be the number of selected CP candidates and $\bm{\tau}^{(d)}=\{{\tau}_{1}^{(d)},\dots,{\tau}_{K^{(d)}}^{(d)}\}$ be the ordered set of CP candidates (${\tau}_{1}^{(d)}<\dots<{\tau}_{K^{(d)}}^{(d)}$ and ${\tau}_{0}^{(d)}=0,{\tau}_{K^{(d)}+1}^{(d)}=T$) for frequency $d\in\{0,\ldots,D-1\}$. Furthermore, we define the set of CP candidate locations as
@@
The first term of the obejective function in (5) indicates the cost for quantifying the variability of segments between two adjacent CP candidates, where $\mathcal{C}\left(\bm{f}_{s:e}^{(d)}\right)$ indicates the cost function for an segment $\bm{f}_{s:e}^{(d)}=(f_{s}^{(d)},f_{s+1}^{(d)},\ldots,f_{e}^{(d)})^{\top}$ specifically defined as
@@
$c_{\text{sym}}^{(d)}=\begin{cases}1&(d=0,\frac{M}{2})\\ 2&(d\neq 0,\frac{M}{2})\end{cases},$
because the frequency spectra have complex conjugate symmetry. The second term indicates the penalty term for the number of CP candidates in each frequency, while the third term indicate the penalty for the number of total CP candidate locations where $(\beta^{(0)},\dots,\beta^{(D-1)})\in\mathbb{R}^{D}$, and $\gamma\in\mathbb{R}$ are hyper-parameters for controlling the balance between the three terms (the details of how to determine these hyper-parameters are presented in Appendix A). The third penalty term indicates that we take into account the trade-off between cost and penalty term for not only $K^{(d)}$ but also $K$ in (5), hence CP candidates of multiple frequencies tend to be detected at the same time point.
@@ 3.2 Approximately Optimizing the Objective Function by Simulated Annealing
Unfortunately, since minimizing the objective function in (5) is a challenging combinatorial optimization problem, we must rely on approximate solutions derived from heuristic algorithms. Following the approach in Lavielle (1998), we employ simulated annealing (Kirkpatrick et al., 1983; Černỳ, 1985) to approximately solve the combinatorial optimization problem. Simulated annealing is a meta-heuristic algorithm widely applied to various practical problems, as it converges asymptotically to a global solution with high probability under specific conditions.
To approximately optimize the objective function in (5), we perform the following two steps:
  • Step 1: Individually estimate the CP candidates for each frequency. Specifically, we solve the problem in (5) with $\gamma=0$. This can be optimally achieved by applying dynamic programming to each sequence.
@@
3.3 Step 1: Generating Initial Solution by Dynamic Programming
@@
3.4 Step 2: Refining Solution by Simulated Annealing
We perform multivariate CP candidate selection in the frequency domain using simulated annealing that is applied for solving large combinatorial optimization problems for which finding global optima is difficult. In each step of simulated annealing, the Metropolis algorithm is used to accept transitions not only to improving solutions that decrease the objective function, i.e., $\Delta E(\mathcal{T}^{\prime},\mathcal{T},\bm{x})=E(\mathcal{T}^{\prime},\bm{x})-E(\mathcal{T},\bm{x})\leq 0$, where $\mathcal{T}$ and $\mathcal{T}^{\prime}$ are current and new solutions, respectively, but also to deteriorating solutions that increase the objective function, i.e., $\Delta E(\mathcal{T}^{\prime},\mathcal{T},\bm{x})>0$, with the probability controlled by a temperature parameter $c$, which allows escape from the local solution. The pseudo code of the Metropolis algorithm is shown in Algorithm 1.

### v2

# Change Point Detection in the Frequency Domain with Statistical Reliability

@@ Abstract
Effective condition monitoring in complex systems requires identifying change points (CPs) in the frequency domain, as the structural changes often arise across multiple frequencies. This paper extends recent advancements in statistically significant CP detection, based on Selective Inference (SI), to the frequency domain. The proposed SI method quantifies the statistical significance of detected CPs in the frequency domain using $p$-values, ensuring that the detected changes reflect genuine structural shifts in the target system. We address two major technical challenges to achieve this. First, we extend the existing SI framework to the frequency domain by appropriately utilizing the properties of discrete Fourier transform (DFT). Second, we develop an SI method that provides valid $p$-values for CPs where changes occur across multiple frequencies. Experimental results demonstrate that the proposed method reliably identifies genuine CPs with strong statistical guarantees, enabling more accurate root-cause analysis in the frequency domain of complex systems.
@@ 1 Introduction
To detect failures in complex systems, it is crucial to accurately identify the frequencies where the significant changes occur. By pinpointing these specific frequencies, the root causes of faults in the systems can be identified. In this paper, we address the change point (CP) detection problem in the frequency domain and propose a method that can identify statistically significant changes in specific frequencies. Quantifying statistical significance in CP detection ensures that the detected changes in the frequency domain reflect genuine structural alterations rather than random noise. Measures such as $p$-values provide a quantitative framework to differentiate real changes from spurious ones, effectively mitigating the risk of incorrect decisions in data-driven systems.
@@
One of the difficulties in frequency-domain CP detection is that changes often appear across multiple frequencies due to shared underlying phenomena influencing broad signal characteristics. Figure 1 shows an example of frequency-domain CP detection problems along with the statistically significant CPs identified using our proposed SI method. Panel (a) shows a time series transformed into signals across five frequencies. Panel (b) presents the CP detection results for each individual frequency $d\in\{d_{0},d_{1},d_{2},d_{3},d_{4}\}$ 111 Note that while the “amplitude” spectra are presented in the figures of this paper to visualize the temporal variations of spectral sequences and the means for each segment, the “complex” spectra are utilized in the actual CP detection and the hypothesis testing (see Section 2 and the subsequent sections for details). . Panel (c) integrates CPs across multiple frequencies, detecting three changes A, B, and C (at CP A, changes occur in frequencies $d_{1}$ and $d_{2}$; at CP B, in $d_{2}$ and $d_{3}$; and at CP C, only in $d_{0}$). In table (d), the obtained $p$-values for changes A, B, and C using the proposed method are 0.008, 0.006, and 0.752 (denoted as selective $p$-values), respectively, indicating that at a significance level of $\frac{0.05}{3}\approx 0.0167$ decided by Bonferroni correction, changes A and B are statistically significant CPs222 In Figure 1, the values labeled as naive $p$-values are inappropriate because they fail to account for double-dipping, leading to excessively small values and an inflated rate of false positive findings. .
To perform statistically significant CP detection, as illustrated in Figure 1, two key technical challenges must be addressed. The first challenge is to extend the SI framework to the frequency domain. We tackle this by formulating the test statistic and the conditioning on the hypothesis selection that accurately account for the properties of discrete Fourier transform (DFT). The second challenge involves quantifying the statistical significance of CPs shared across multiple frequencies. Addressing this requires solving a combinatorial optimization problem, which is computationally infeasible to solve globally optimally, necessitating a heuristic approach to obtain the approximate solution. In this study, we employ a method based on simulated annealing (SA) (Kirkpatrick et al., 1983; Černỳ, 1985) and implement the SI framework to appropriately quantify the statistical significance of the approximate solution derived from the heuristic algorithm.
The rest of this paper is organized as follows. Section 2 formulates the problem. Section 3 explains the heuristic algorithm used to select CP candidates in the first stage of the proposed method. Section 4 describes the method for quantifying the statistical significance of CP candidates using the SI framework in the second stage. In Section 5, we demonstrate the effectiveness of our proposed method through comprehensive numerical experiments using both synthetic and real-world data. For reproducibility, our implementation is available at https://github.com/Takeuchi-Lab-SI-Group/si_for_frequency-domain_change_point_detection. Finally, Section 6 concludes this paper.
@@
Our contribution.
The main contribution of this study is the development of a statistical testing method, based on the framework of SI, to quantify the statistical significance of detected CPs in the frequency domain. While CP detection in the frequency domain is an important problem in engineering and several algorithms have been proposed, to the best of our knowledge, no existing method properly quantifies the statistical significance of CPs in the form of $p$-values. By leveraging the recently advancing SI framework, we provide a solution to this unresolved issue. There are various CP detection algorithm in the frequency domain. In this study, as an example, we consider a CP detection algorithm that identifies changes common across multiple frequencies using dynamic programming (DP) and simulated annealing. It should be noted that our contribution does not lie in this CP detection algorithm itself, but in proposing a statistical test to quantify the significance of the detected CPs.
Related work.
The CP detection problem has long been studied with various applications in a variety of fields, such as finance (Fryzlewicz & Subba Rao, 2014; Pepelyshev & Polunchenko, 2017), bioinformatics (Chen & Wang, 2008; Muggeo & Adelfio, 2011; Pierre-Jean et al., 2015), climatology (Reeves et al., 2007; Beaulieu et al., 2012), and machine monitoring (Lu et al., 2017; 2018). In the statistics and machine learning communities, various methods have been proposed for identifying multiple CPs from univariate sequences. The most straightforward approach is repeatedly applying single CP detection algorithms, such as binary segmentation (Scott & Knott, 1974). Examples of such approaches include circular binary segmentation (Olshen et al., 2004) and wild binary segmentation (Fryzlewicz, 2014). Another line of research has proposed numerous approaches based on penalized likelihood, such as Segment Neighbourhood (Auger & Lawrence, 1989), Optimal Partitioning (Jackson et al., 2005), PELT (Killick et al., 2012), and FPOP (Maidstone et al., 2017). In these studies, dynamic programming is employed to solve the penalized likelihood minimization problem. While most CP detection studies focus on the time domain, a few studies have targeted the frequency domain, such as Adak (1998), Last & Shumway (2008), and Preuss et al. (2015).
@@
SI was initially introduced as a method of statistical inference for feature selection in linear models (Taylor & Tibshirani, 2015; Fithian et al., 2015) and later extended to various feature selection algorithms, including marginal screening (Lee & Taylor, 2014), stepwise feature selection (Tibshirani et al., 2016), and Lasso (Lee et al., 2016). The core concept of SI is to derive the exact null distribution of the test statistic conditional on the hypothesis selection event, thereby enabling valid statistical inference with controlled type I error rate. Recently, significant attention has been given to applying SI to more complex supervised learning algorithms, such as kernel models (Yamada et al., 2018), boosting (Rügamer & Greven, 2020), tree-structured models (Neufeld et al., 2022), and neural networks (Duy et al., 2022; Miwa et al., 2023; Shiraishi et al., 2024b). Furthermore, SI has proven valuable for unsupervised learning tasks, including clustering (Lee et al., 2015; Chen & Witten, 2023; Gao et al., 2024), outlier detection (Chen & Bien, 2020; Tsukurimichi et al., 2022), domain adaptation (Duy et al., 2024), and segmentation (Tanizaki et al., 2020; Duy et al., 2022). SI was first used for CP detection problem in Hyun et al. (2018), which focused on Fused Lasso algorithm. Since then, the framework has been explored in various CP detection algorithms, including the CUSUM-based method (Umezu & Takeuchi, 2017), binary segmentation and its variants (Hyun et al., 2021), dynamic programming (Duy et al., 2020), and other related problems (Sugiyama et al., 2021; Jewell et al., 2022; Carrington & Fearnhead, 2024; Shiraishi et al., 2024a). Existing studies on CP detection have focused on the time domain, making this the first to offer valid statistical inferences for frequency-domain CPs using the SI framework.
@@ 2.2 Statistically Significant CP Detection in the Frequency Domain
As discussed in Section 1, we have prior knowledge that simultaneous changes occur across multiple distinct frequencies. In Section 3, we introduce a heuristic algorithm that generates CP candidates by incorporating this prior knowledge. Subsequently, in Section 4, we present an SI framework to quantify the statistical significance of each CP candidate in the form of $p$-values. Finally, we select the candidates with $p$-values smaller than the user-specified significance level (e.g., 0.05 or 0.01) as our final CPs.
@@ 3 Heuristic Algorithm for CP Candidate Selection in Multiple Frequencies
As discussed in Section 1, the types of problems occurring in time series signals can be systematically determined by simultaneously detecting CPs across multiple frequencies, such as the harmonics and sidebands of the characteristic frequencies. Thus, we formulate CP candidate selection as an optimization problem that aims to not only estimate the optimal number and location of CPs for each frequency, but also reduce the total number of change locations across all frequencies by aligning the positions.
@@ 3.1 Objective Function for CP Candidate Selection
Let $K^{(d)}$ be the number of selected CP candidates and $\bm{\tau}^{(d)}=\{{\tau}_{1}^{(d)},\dots,{\tau}_{K^{(d)}}^{(d)}\}\subseteq[T-1]$ be the ordered set of CP candidates (${\tau}_{1}^{(d)}<\dots<{\tau}_{K^{(d)}}^{(d)}$ and ${\tau}_{0}^{(d)}=0,{\tau}_{K^{(d)}+1}^{(d)}=T$) for frequency $d\in\{0,\ldots,D-1\}$. Furthermore, we define the set of CP candidate locations as
@@
The first term of the obejective function in (5) indicates the cost for quantifying the variability of segments between two adjacent CP candidates, where $\mathcal{C}\left(\bm{f}_{s:e}^{(d)}\right)$ indicates the cost function for a segment $\bm{f}_{s:e}^{(d)}=(f_{s}^{(d)},f_{s+1}^{(d)},\ldots,f_{e}^{(d)})^{\top}$ specifically defined as
@@
$c_{\text{sym}}^{(d)}=\begin{cases}1&\text{{if $d=0,\frac{M}{2}$}},\\ 2&\text{{if $d\neq 0,\frac{M}{2}$}},\end{cases}$ (6)
because the frequency spectra have complex conjugate symmetry. The second term indicates the penalty term for the number of CP candidates in each frequency, while the third term indicate the penalty for the number of total CP candidate locations, where $(\beta^{(0)},\dots,\beta^{(D-1)})\in\mathbb{R}^{D}$ and $\gamma\in\mathbb{R}$ are hyper-parameters for controlling the balance between the three terms (the details of how to determine these hyper-parameters are presented in Appendix A). The third penalty term indicates that we take into account the trade-off between cost and penalty term for not only $K^{(d)}$ but also $K$ in (5), hence CP candidates of multiple frequencies tend to be detected at the same time point.
@@ 3.2 Approximately Optimizing the Objective Function by Simulated Annealing
Unfortunately, since minimizing the objective function in (5) is a challenging combinatorial optimization problem, we must rely on approximate solutions derived from heuristic algorithms. Following the approach in Lavielle (1998), we employ simulated annealing (Kirkpatrick et al., 1983; Černỳ, 1985) to approximately solve the optimization problem. Simulated annealing is a meta-heuristic algorithm widely applied to various practical problems, as it converges asymptotically to a global solution with high probability under specific conditions. To approximately optimize the objective function in (5), we perform the following two steps:
  • Step 1: Individually estimate the CP candidates for each frequency. Specifically, we minimize the objective function in (5) with $\gamma=0$. This can be optimally achieved by applying dynamic programming to each sequence.
@@
The reason for obtaining approximate solutions using simulated annealing is that the global minimization of the objective function in (5) across multiple frequencies cannot be achieved by directly extending the dynamic programming approach used in step 1. This inability to obtain the global optimum stems from the fact that it is impossible to construct a Bellman equation for this optimization problem; specifically, additional information which is not preserved in the solutions of subproblems may be needed at later-stage decisions.
@@
3.3 Step 1: Generating an Initial Solution by Dynamic Programming
@@
3.4 Step 2: Refining the Solution by Simulated Annealing
We perform multivariate CP candidate selection in the frequency domain using simulated annealing that is applied for solving large combinatorial optimization problems for which finding global optima is difficult. In each step of simulated annealing, the Metropolis algorithm is used to accept transitions not only to improving solutions that decrease the objective function, i.e., $\Delta E(\bm{\mathcal{T}}^{\prime},\bm{\mathcal{T}},\bm{x})=E(\bm{\mathcal{T}}^{\prime},\bm{x})-E(\bm{\mathcal{T}},\bm{x})\leq 0$, where $\bm{\mathcal{T}}$ and $\bm{\mathcal{T}}^{\prime}$ are current and new solutions, respectively, but also to deteriorating solutions that increase the objective function, i.e., $\Delta E(\bm{\mathcal{T}}^{\prime},\bm{\mathcal{T}},\bm{x})>0$, with the probability controlled by a temperature parameter $c$, which allows escape from the local solution. The pseudo code of the Metropolis algorithm is shown in Algorithm 1.